# leftmargin=0pt, rightmargin=0pt, vskip=0pt

### v1

@@ Abstract.
Understanding sensor data can be challenging for non-experts because of the complexity and unique semantic meanings of sensor modalities. This calls for intuitive and effective methods to present sensor information. However, creating intuitive sensor data visualizations presents three key challenges: the variability of sensor readings, gaps in domain comprehension, and the dynamic nature of sensor data. To address these issues, we develop Vivar, a novel AR system that integrates multi-modal sensor data and presents 3D volumetric content for visualization. In particular, we introduce a cross-modal embedding approach that maps sensor data into a pre-trained visual embedding space through barycentric interpolation. This allows for accurate and continuous integration of multi-modal sensor information. Vivar also incorporates sensor-aware AR scene generation using foundation models and 3D Gaussian Splatting (3DGS) without requiring domain expertise. In addition, Vivar leverages latent reuse and caching strategies to accelerate 2D and AR content generation. Our extensive experiments demonstrate that our system achieves 11$\times$ latency reduction without compromising quality. A user study involving over 485 participants, including domain experts, demonstrates Vivar’s effectiveness in accuracy, consistency, and real-world applicability, paving the way for more intuitive sensor data visualization.
@@ 1   Introduction
Although sensor data offers immense utility, its inherent complexity and abstract nature present significant challenges, particularly for non-experts seeking to interpret it. Raw data often contain intricate relationships, nuanced variations, and diverse modalities, making it difficult to derive meaningful insights at a glance. For example, a Perfusion Index reading of $3.5\%$, indicating moderate blood flow, may be clear to clinicians but confusing for non-experts unfamiliar with its significance for health monitoring (Elshal et al., 2021). To address this gap, intuitive and accessible visualization methods are essential. These methods can enable users to transform abstract data into interpretable forms that facilitate better decision-making and situational awareness. In doing so, they can also expand the adoption of sensor-driven technologies.
@@
Augmented Reality (AR) offers a particularly compelling medium for sensor information visualization to bridge this domain gap. By overlaying digital content onto the physical world, AR provides an immersive and context-aware experience that enhances user interaction with data (Malkawi and Srinivasan, 2005; Reipschläger et al., 2020; Hou et al., 2024). This capability allows users to intuitively engage with sensor information in their surroundings, turning complex, multimodal data into tangible, interactive experiences. For instance, AR can overlay environmental information such as air quality (An et al., 2024), temperature (Wilson et al., 2023), or safety conditions (Gong et al., 2024) as the representative manifestations onto physical spaces, offering a clear and accessible representation of these parameters without requiring specialized expertise. The immersive nature of AR effectively lowers barriers to data interpretation, benefiting both experts and non-experts, as shown in Figure 1.
@@
Presenting sensor information in AR to bridge the domain gap introduces three major challenges. 1) Variability of Sensor Readings: Sensor measurements often vary significantly, both individually and in combination. For example, temperature readings may range from -30°C to 40°C, representing vastly different conditions (Mahendran et al., 2024). Additionally, sensors frequently work together, measuring variables like air quality, temperature, and pollution levels simultaneously (Bakirci, 2024), making it challenging to coherently blend these interrelated data points into a unified scene. 2) Domain Gap in Comprehension: The abstract nature of sensor data creates a significant gap in human understanding, requiring tailored visual representations to effectively communicate key features. Bridging this gap often involves long-term study and expert interpretation. 3) Dynamic Nature of Sensor Data: The ever-changing nature of sensor readings necessitates just-in-time visualizations that match the rate of data updates, requiring systems capable of rapid interpretation and immediate visual updates.
@@
Existing sensor data visualization methods face substantial limitations. Traditional dashboards, while effective for domain experts, often prove inaccessible to non-technical users. Recent advances in AI-driven visualization, such as Artificial Intelligence-Generated Content (AIGC) frameworks like Stable Diffusion (Esser et al., 2024), struggle to seamlessly integrate sensor data as inputs. Similarly, three-dimensional generation methods (Tang et al., 2023; Poole et al., 2022) show promise but are hindered by high computational demands and limited capability for generating volumetric sensor presentation. These shortcomings highlight the need for a new solution that combines intuitive and immersive user experiences with efficient performance.
@@
To address these challenges, we propose Vivar, a novel system for intuitive and efficient sensor data visualization in AR. Central to Vivar is its use of barycentric interpolation (Grünbaum et al., 1967; Angel, 2005), a technique inspired by shading principles, to map diverse sensor data into a shared visual embedding space. This interpolation ensures smooth transitions and accurate blending of sensor readings, enabling seamless visualization in both single-sensor and multi-sensor scenarios. The system features an automated, sensor-aware pipeline that generates AR scenes directly from sensor readings, bridging the gap between abstract data and intuitive representation. To manage the rapid dynamics of sensor data, Vivar incorporates a latent-reuse mechanism, caching 2D and AR content for timely responsiveness and enhanced efficiency.
@@
    Just-in-time efficiency: We create a fast generation framework with latent-reuse and caching strategies for 2D and AR content that speed the generation up to 10x compared to existing models.
@@ 2   Related Work
Presenting the sensor data information to the users has been studied in both industry and academia, with tools like SensorViz (Kim et al., 2022) and immersive technologies for environmental data (McGee et al., 2024; Chun et al., 2023) providing detailed representations of specific sensor types in industries, smart cities, and traffic systems. However, these tools often cater to domain experts and fail to bridge the gap for non-expert users who struggle with interpreting complex and abstract sensor data.
Presenting sensor data within its real-world context is critical. Systems like DeepSee (Coscia et al., 2024), Virtual Dream Reliving (Liu et al., 2024), and ImmersiveFlora (Hu et al., 2024) overlay raw sensor data onto physical environments, improving interpretability in fields like environmental monitoring and healthcare. Other approaches use robots, avatars, or narrative techniques (Schömbs et al., 2024; Sareen et al., 2019; Yonezawa et al., 2001; Reipschläger et al., 2022; Bhatia et al., 2024), yet they often lack seamless transitions across data modalities.
As access to sensor data grows, addressing these challenges is increasingly urgent. Effective visualizations enhance comprehension (Kurze et al., 2020; Quadri et al., 2024; Nobre et al., 2024), with methods like metaphorical mappings or prompt-based teaching (Tkachev et al., 2022; Almeda et al., 2024) showing promise. Yet, existing models are limited in handling heterogeneous sensor modalities. A system capable of intuitive, contextually relevant representations is essential for empowering both experts and the general public.
@@
Generative models such as Stable Diffusion (Rombach et al., 2022) and DreamFusion (Poole et al., 2022) have gained prominence for their ability to create synthetic content and visualize complex data. These models excel in generating realistic imagery and 3D structures, with applications in domains ranging from art to scientific visualization. While they demonstrate significant potential in bridging abstract data and tangible representations, their integration with environments with dynamic sensor readings remains underexplored. Attempts to incorporate generative models into AR systems focus on creating immersive and adaptable 3D content. For instance, methods like SharedNeRF (Sakashita et al., 2024) enable volumetric reconstructions for AR applications, offering realistic representations but facing challenges with computational overhead and latency. Models like 3DGS (Kerbl et al., 2023) have emerged to address these issues, improving efficiency for real-time applications. However, adapting these systems for sensor data visualization presents unique difficulties, such as maintaining semantic coherence while dynamically representing multimodal inputs (Li et al., 2020; Zhang et al., 2023c; Chen et al., 2024; Lin et al., 2023; Fu et al., 2024).
@@
3   Motivation
@@
Figure 2. Cosine similarity versus numerical differences using different multimodal encoders: CLIP, VisualBERT, and ALIGN. The figures illustrate (a) biases and (b-c) the inability to effectively embed sensor data.
Sensors play a crucial role in capturing and providing valuable insights across domains such as healthcare, environmental monitoring, and industrial automation. However, the complexity and abstract nature of sensor data often make it inaccessible to non-experts, limiting its broader adoption and practical utility. Current visualization tools, such as dashboards, primarily target domain experts, leaving non-expert users struggling to interpret fragmented data or derive meaningful insights from diverse modalities. Furthermore, numerical readings, while precise, fail to intuitively convey actionable information. For example, a pressure reading of “750 hPa” is technically accurate but lacks immediate relevance for users unfamiliar with its implications for altitude or weather conditions.
Recent advancements in embedding approaches provide promising solutions by transforming sensor data into compact and meaningful representations. Unified embedding spaces enable seamless integration of sensor data with other modalities, such as text or images. However, current methods face several challenges, including inadequate alignment with sensor semantics (Ma et al., 2024; Kim et al., 2024) and biases in data interpretations (Meng et al., 2024). State-of-the-art multimodal encoders, such as CLIP (Radford et al., 2021), VisualBERT (Li et al., 2019), and ALIGN (Jia et al., 2021), struggle to map numerical sensor values into meaningful embeddings effectively. These shortcomings are evident in the inconsistencies between embedding correlations and the actual numerical relationships in sensor data (Figure 2). In this experiment, we embed sensor readings in the form of sentences such as ”The temperature is $x$ Celsius.” into the embedding space with the above encoders and calculate their cosine similarity of embeddings in pairs relative to the numerical differences of the sensor readings. We average the data point in blue vertically to obtain the trend shown in orange. Ideally, if the embeddings accurately represent the numerical values, the relative relationships between numbers would be preserved. Specifically, larger numerical differences should correspond to lower cosine similarity in the embedding space. The results, shown in the figure, highlight several key issues: (a) Biases in embedding numerical readings: While the overall trend decreases, the embeddings exhibit proximity to others at similar digits, reflecting periodic patterns that do not align with actual numerical relationships. (b) Inability to capture semantic changes: The models fail to adequately reflect changes in sensor semantics, resulting in embeddings that are inconsistent with the underlying physical dynamics of the data.
To address these gaps, we propose Vivar, a novel system that leverages cross-modal embeddings and sensor-aware generation to deliver seamless, accurate, and immersive AR visualizations of sensor data.
@@ 4   Design Overview
Providing accurate and intuitive sensor information visualization presents three main challenges: 1) Multi-Modality: Sensor readings often originate from diverse sources and formats, exhibiting significant variability both individually and in combination, which makes seamless integration and interpretation challenging. 2) Context and Domain Differences: The meaning of sensor data varies across different contexts, requiring tailored visual representations and domain expertise for accurate interpretation. This domain gap often demands substantial effort to bridge effectively. 3) Temporal Dynamics: Sensor data frequently change over time, necessitating visualizations that can adapt and update timely to reflect dynamic data accurately.
To address these challenges, we present Vivar, a system designed to transform abstract sensor data into intuitive and immersive AR experiences. Our system collects and integrates sensor data from various sources onto a centralized server. For example, AQI and climate information are sourced from weather stations, while home sensor data is obtained via APIs such as Apple HomeKit (Inc., 2024). The collected data is then processed on the server to blend and generate AR scenes. Finally, users can view the information through the AR glasses.
Vivar leverages cross-modal embeddings and sensor-aware visualization techniques to provide actionable, context-rich representations. As shown in Figure 3, the system operates in two key phases:
  • Barycentric Interpolation for Cross-modal Sensor Embedding: We propose a unified approach to embed different types of sensor readings into a shared embedding space. This method first embeds the representative sensor readings into the embedding space and uses them as anchors to connect numerical readings to points in the embedding space. Then, we design the embedding interpolation using barycentric coordinates, inspired by shading principles, to calculate embeddings for new readings. This approach ensures smooth transitions between sensor embeddings while maintaining a consistent and interpretable representation for both single and multiple sensor types.
  • Sensor-Aware Visual Production: To provide intuitive and domain-enriched visualization of sensor readings with fast response times, we designed a sensor-aware visual production framework. This framework automatically integrates sensor context information with their schemas and manifestations to effectively represent sensor data. Using a customized volumetric 3D production model, we integrate spatial density of sensor information into the visualization. By implementing latent reuse caches at each production layer, we achieve response times that are 10$\times$ faster than baseline 3D generation pipelines.
@@ 5.1. CLIP as the Sensor Visual Space
Despite its strengths, CLIP is primarily designed for text and image data, whereas sensor readings are numerical and present unique integration challenges. Mapping sensor information to visualization content requires a carefully designed function that translates raw data into meaningful visual representations within the shared “visual space.” The key objective is constructing a mapping function that accurately positions sensor data within this embedding space.
@@
Previous studies have shown that interpolating between two embeddings enables image generation models, such as Stable Diffusion (Face, 2023), to create smooth transformations from one generated image to another (Face, 2023; Podell et al., 2023). Inspired by this, we propose an anchor-based interpolation method to generate embeddings for new sensor readings by interpolating between anchor embeddings.
@@
The embeddings for these anchors are computed from the text embeddings of the sensor readings using a pre-trained CLIP text embedding model. Subsequently, these anchor embeddings are blended to generate embeddings for new sensor readings. As illustrated in Figure 5(a), for a single sensor reading, such as temperature, the embedding can be derived through linear interpolation between the corresponding anchors. For scenarios involving more than two anchors (e.g., with multiple representative points), the embedding is calculated using linear interpolation between the neighboring anchors (Figure 5(b)).
Key challenges arise from the increasing number of sensors and anchors and the complexity of handling their combinations:
@@
We present a showcase example with humidity sensors to demonstrate the effectiveness of our method in visual space. We generate images using Stable Diffusion with embeddings computed by CLIP, based on the prompt template: ”A leaf representing $x\%$ relative humidity”, where $x\in\{10,22,41,71,89\}$. We compare the results of our embedding interpolation method (referred to as the after set) with a baseline approach (referred to as the before set), where embeddings were directly fed into the diffusion model. The comparative results are shown in Figure 6. For interpolation, we use embeddings generated at the minimum and maximum values of $x$. In both sets, the first and last images are identical. However, differences emerge in the intermediate images.
@@
Ideally, as humidity increases, the generated images should depict leaves that progressively appear healthier and greener, exhibiting less withering. While this progression is evident in the after set, it is inconsistent in the before set. For instance, the image corresponding to 71% relative humidity (the fourth image from the left) in the before set fails to align with the expected progression.
To quantify these differences, we computed the cosine similarity of each embedding relative to the embeddings of the two extreme conditions (the driest and most humid cases). This analysis yielded two sequences from the before set, shown as blue lines in the top two plots in Figure 6. The corresponding sequences from the after set are shown as orange lines in the same plots. To further evaluate the disruption in monotonic progression, we employed Kendall’s Tau coefficient, which measures the degree of monotonicity, with a value approaching 1 indicates a perfect monotonic increase, while a value approaching -1 indicates a perfect monotonic decrease. It is defined as $\tau=\frac{2(P-Q)}{n(n-1)}$, where $P$ is the number of concordant pairs, $Q$ is the number of discordant pairs, and $n$ is the total number of pairs.
The results indicate that the embedding interpolation method achieves perfect monotonicity, as reflected by Kendall’s Tau values and the smooth progression of the blue curves. This contrasts with the inconsistencies observed in the before set, as evident in the valley of the blue curve.
This example demonstrates that the embedding interpolation approach provides a more consistent and accurate representation of sensor data in the sensor visual space. By interpolation embeddings from anchor points, we achieve smooth transitions in the generated images corresponding to incremental changes in sensor readings. This not only enhances visual coherence but also ensures that the embeddings accurately reflect the underlying sensor semantics. Next, we extend this design to cases with more than one sensor reading.

### v2

@@ Abstract.
Understanding sensor data can be difficult for non-experts because of the complexity and different semantic meanings of sensor modalities. This leads to a need for intuitive and effective methods to present sensor information. However, creating intuitive sensor data visualizations presents three key challenges: the variability of sensor readings, gaps in domain comprehension, and the dynamic nature of sensor data. To address these issues, we propose Vivar, a novel system that integrates multi-modal sensor data and presents 3D volumetric content for AR visualization. In particular, we introduce a cross-modal embedding approach that maps sensor data into a pre-trained visual embedding space through barycentric interpolation. This approach accurately reflects value changes in multi-modal sensor information, ensuring that sensor variations are properly shown in visualization outcomes. Vivar also incorporates sensor-aware AR scene generation using foundation models and 3D Gaussian Splatting (3DGS) without requiring domain expertise. In addition, Vivar leverages latent reuse and caching strategies to accelerate 2D and AR content generation, demonstrating $\times1$ latency reduction without compromising quality. A user study involving over 503 participants, including domain experts, demonstrates Vivar’s effectiveness in accuracy, consistency, and real-world applicability, paving the way for more intuitive sensor data visualization.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1   Introduction
Although sensor data offers immense utility, its inherent complexity and abstract nature present significant challenges, particularly for non-experts seeking to interpret it. Raw data often contain intricate relationships, nuanced variations, and diverse modalities, making it difficult to derive meaningful insights at a glance. For example, an Air Quality Index (AQI) reading of 157, which indicates unhealthy air conditions, may be readily understood by environmental professionals but could confuse non-specialists who are less familiar with its implications for public health and outdoor activities. To address this gap, intuitive and accessible visualization methods are essential, enabling transforming abstract data into interpretable forms that help better decision-making, situational awareness, and broader adoption of sensor-driven technologies.
@@
Augmented Reality (AR) offers a particularly compelling medium for sensor information visualization to bridge this domain gap. By overlaying digital content onto the physical world, AR provides an immersive and context-aware experience that enhances user interaction with data (Malkawi and Srinivasan, 2005; Reipschläger et al., 2020). This capability allows users to intuitively engage with sensor information in their surroundings, turning complex, multimodal data into tangible, interactive experiences. For instance, AR can overlay environmental information such as air quality (An et al., 2024), temperature (Wilson et al., 2023), safety conditions (Gong et al., 2024), or military training scenarios where the U.S. Air Force leverages multi-spectral operational mixed reality training devices to introduce new training efficiencies while overcoming previous technology limitations (Technology, 2023) as the representative manifestations onto physical spaces, offering a clear and accessible representation of these parameters without requiring specialized expertise. The immersive nature of AR effectively lowers barriers to data interpretation, benefiting both experts and non-experts, as shown in Figure 1.
@@
Visualizing sensor information in AR to bridge the gap introduces three major challenges. 1) Variability of Sensor Readings: Sensor measurements often vary significantly, both individually and in combination. For example, temperature readings may range from -30°C to 40°C, representing vastly different conditions (Mahendran et al., 2024). Additionally, sensors frequently work together, measuring variables like air quality, temperature, and pollution levels simultaneously (Bakirci, 2024), making it challenging to coherently blend these interrelated data points into a unified scene. 2) Domain Gap in Comprehension: The abstract nature of sensor data creates a significant gap in human understanding, requiring tailored visual representations to effectively communicate key features. Bridging this gap often involves long-term study and expert interpretation. 3) Dynamic Nature of Sensor Data: The ever-changing nature of sensor readings necessitates just-in-time visualizations that match the rate of data updates, requiring systems capable of rapid interpretation and immediate visual updates.
@@
Existing sensor data visualization methods face substantial limitations. Traditional dashboards, while effective for domain experts, often prove inaccessible to non-technical users. Recent advances in AI-driven visualization, such as Artificial Intelligence Generated Content (AIGC) frameworks like Stable Diffusion (Esser et al., 2024), struggle to seamlessly integrate sensor data as inputs. Similarly, three-dimensional generation methods (Tang et al., 2023; Poole et al., 2022; Zhou et al., 2025) show promise but are hindered by high computational demands and limited capability for generating volumetric sensor presentation. These shortcomings highlight the need for a new solution that combines intuitive and immersive user experiences with efficient performance.
@@
To address these challenges, we propose Vivar, a novel system for intuitive and efficient sensor data visualization in AR. Central to Vivar is its use of barycentric interpolation (Grünbaum et al., 1967; Angel, 2005), a technique inspired by shading principles, to map diverse sensor data into a shared visual embedding space. This interpolation ensures smooth transitions and accurate blending of sensor readings, enabling seamless visualization in both single-sensor and multi-sensor scenarios. The system features an automated, sensor-aware pipeline that generates AR scenes directly from sensor readings, bridging the gap between abstract data and intuitive representation. To manage the rapid dynamics of sensor data, Vivar incorporates a latent-reuse mechanism, caching 2D and AR content for timely responsiveness and enhanced efficiency. We have implemented Vivar as an end-to-end system, encompassing data ingestion, visualization generation, and AR rendering. We will open-source our implementation to contribute to the field and inspire broader applications.
@@
    Just-in-time efficiency: We create a fast generation framework with latent-reuse and caching strategies for 2D and AR content that speed the generation up to $\times1$ compared to existing models.
@@ 2   Related Work
Presenting the sensor data information to the users has been studied in both industry and academia, with tools like SensorViz (Kim et al., 2022) and immersive technologies for environmental data (McGee et al., 2024; Chun et al., 2023) providing detailed representations of specific sensor types in industries, smart cities, and traffic systems. However, these tools often cater to domain experts and fail to benefit non-expert users who struggle with interpreting complex and abstract sensor data.
Systems like DeepSee (Coscia et al., 2024), Virtual Dream Reliving (Liu et al., 2024), and ImmersiveFlora (Hu et al., 2024) overlay raw sensor data onto physical environments, improving interpretability in fields like environmental monitoring and healthcare. Other approaches use robots, avatars, or narrative techniques (Schömbs et al., 2024; Sareen et al., 2019; Yonezawa et al., 2001; Reipschläger et al., 2022; Bhatia et al., 2024), yet they often lack seamless transitions across data modalities.
As access to sensor data grows, addressing these challenges is increasingly urgent. Effective visualizations enhance comprehension (Kurze et al., 2020; Quadri et al., 2024; Nobre et al., 2024; Rinaldi et al., 2024; Chen et al., 2024a), with methods like metaphorical mappings or prompt-based teaching (Tkachev et al., 2022; Almeda et al., 2024) showing promise. However, existing models are limited in handling heterogeneous sensor modalities. A system capable of intuitive, contextually relevant representations is essential for empowering both experts and the general public.
@@
Generative models such as Stable Diffusion (Rombach et al., 2022) and DreamFusion (Poole et al., 2022) have gained prominence for their ability to create synthetic content and visualize complex data. These models excel in generating realistic imagery and 3D structures, with applications in domains ranging from art to scientific visualization (Zhu et al., 2024). While they demonstrate significant potential in bridging abstract data and tangible representations, their integration with environments with dynamic sensor readings remains underexplored. Attempts to incorporate generative models into AR systems focus on creating immersive and adaptable 3D content. For instance, methods like SharedNeRF (Sakashita et al., 2024) enable volumetric reconstructions for AR applications, offering realistic representations but facing challenges with computational overhead and latency. Models like 3DGS (Kerbl et al., 2023) have emerged to address these issues, improving efficiency for real-time applications. However, adapting these systems for sensor data visualization presents unique difficulties, such as maintaining semantic coherence while dynamically representing multimodal inputs (Li et al., 2020; Zhang et al., 2023c; Chen et al., 2024b; Lin et al., 2023).
@@
3   Motivation Study
Challenges in Understanding Sensor Data
To understand how well people interpret sensor data and the necessity for visual representations, we conducted a survey with 122 participants across various age groups and technical backgrounds. Participants were presented with a series of common sensor measurements – including air quality indices (AQI), temperature, humidity, and ultraviolet (UV) index – and asked to interpret these values and their implications for everyday decisions. Additionally, they evaluated alternative representation methods that used visual objects to embody the same data. Following are some examples in the survey:
  1. (1)
    Air Quality Index: Participants were shown the reading “AQI: 157” and asked if outdoor activities were safe. Later, the same condition was presented as a visual scene with visible smog and reduced visibility.
  2. (2)
    Environmental Comfort: Participants were asked to interpret the meaning of “Temperature: 23°C, Humidity: 85%” for human comfort. The same condition was later visualized as a room with visible moisture on surfaces.
  3. (3)
    UV Exposure: Participants interpreted “UV Index: 8” and recommended protective measures. This was later represented visually as intense sunlight causing visible effects on exposed surfaces.
Figure 2. Survey About Sensor Data Interpretation
As show in Figure 2, this survey highlighted notable challenges for people in interpreting numerical sensor data: only 47% of participants correctly identified safe versus unsafe AQI levels, while 38% misinterpreted relative humidity readings. Additionally, 72% of participants reported experiencing medium to high cognitive load when simultaneously integrating multiple sensor readings. However, when presented with object-based visualizations of the same sensor data, comprehension accuracy improved by an average of 34% across all sensor types. The most significant improvement (52%) was observed among non-technical participants. Participants were also asked about their preferences regarding different visualization approaches, specifically comparing AR-based visualization to traditional flat-screen methods. 52% of participants considered AR the most intuitive visualization method for experiencing sensor data, while 33% preferred flat images, and 15% preferred numerical data. These results highlight both the difficulties people face when interpreting sensor data and the potential of intuitive visual representations – especially in AR contexts – to bridge this comprehension gap, particularly for users without domain expertise.
Foundation Model Limitations
@@ 3   Motivation Study
Figure 3. Relationship between cosine similarity and numerical differences for three multimodal encoders. All three models demonstrate cognitive biases or limitations when embedding quantitative data.
The previous survey demonstrates that visualization provides people with a more tangible way to understand abstract sensor data. However, creating effective visualizations presents a significant challenge: there is no comprehensive database that maps arbitrary sensor values to appropriate visual representations across diverse environments and contexts. This gap leads a generative approach that can dynamically create suitable visualizations tailored to specific sensor readings. This approach requires pretrained multimodal models that can convert numerical sensor data into compact and meaningful representations. These models enable the creation of unified embedding spaces where sensor readings can be seamlessly integrated with visual elements, facilitating the generation of contextually appropriate visualizations that enhance human comprehension of complex sensor data.
However, this approach requires models to be sensitive to semantic changes in numerical values. We conducted an experiment to assess how state-of-the-art multimodal encoders – CLIP (Radford et al., 2021), VisualBERT (Li et al., 2019), and ALIGN (Jia et al., 2021) – comprehend numbers in context. We encoded sentences like ”The temperature is $x$ Celsius” into embedding spaces, and calculated the cosine similarity between embeddings relative to the numerical differences, as shown in Figure 3. Each blue point represents: $x=|v_{1}-v_{2}|$, $y=\text{sim}(\mathcal{E}(\mathcal{C}(v_{1})),\mathcal{E}(\mathcal{C}(v_{2})))$ , where $\mathcal{C}(v)$ is the contextual sentence with value $v$, $\mathcal{E}$ is the encoder, and sim is cosine similarity. The orange line shows the vertical average of these points. Ideally, if the embeddings accurately reflect changes in numerical values, the relative relationships between numbers would be preserved. Specifically, larger numerical differences should correspond to lower cosine similarity in the embedding space. The results, shown in the figure, highlight several key issues:
  1. (a)
    Biases in embedding numerical readings: While the overall trend decreases, the embeddings exhibit proximity to others at similar digits, reflecting periodic patterns that do not align with actual numerical relationships.
  2. (b)
    Inability to capture semantic changes: The models fail to adequately reflect changes in sensor semantics, resulting in embeddings that are inconsistent with the underlying physical dynamics of the data.
Addressing these gaps requires a system that can effectively convert sensor data into intuitive and accurate AR visualizations through cross-modal embedding and sensor-aware generation methods.
@@ 4   Design Overview
Providing accurate and intuitive sensor information visualization presents three main challenges: 1) Multi-Modality: Sensor readings often originate from diverse sources and formats, making seamless integration and interpretation challenging. 2) Context and Domain Differences: The meaning of sensor data varies across different contexts, requiring tailored visual representations for accurate interpretation. 3) Temporal Dynamics: Sensor data frequently change over time, necessitating visualizations that adapt to real-time changes.
To address these challenges, we present Vivar, a system transforming abstract sensor data into intuitive and immersive AR experiences. As shown in Figure 4, the system operates in two key phases:
  • Barycentric Interpolation for Cross-modal Sensor Embedding: We embed different types of sensor readings into a shared embedding space by first embedding representative sensor readings as anchors. Our interpolation uses barycentric coordinates, inspired by shading principles, to calculate embeddings for new readings, ensuring smooth transitions while maintaining consistent representation across sensor types.
  • Sensor-Aware Visual Production: Our sensor-aware visual production framework automatically integrates sensor context into visualizations. Using a customized volumetric 3D production model, we integrate spatial density of sensor information. Latent reuse caches at each production layer achieve response times 10$\times$ faster than baseline 3D generation pipelines.
@@ 5.1. CLIP as the Sensor Visual Space
Despite its strengths, CLIP is primarily designed for text and image data, whereas sensor readings are numerical and present unique integration challenges as we have shown in Figure 3. Mapping sensor information to visualization content requires a carefully designed function that translates raw data into meaningful visual representations within the shared “visual space.” The key objective is constructing a mapping function that accurately positions sensor data within this embedding space.
@@
Previous studies (Hou et al., 2024) have shown that interpolating between two embeddings enables image generation models, such as Stable Diffusion (Face, 2023), to create smooth transformations from one generated image to another (Face, 2023; Podell et al., 2023). Inspired by this, we used an anchor-based interpolation method to generate embeddings for new sensor readings by interpolating between anchor embeddings.
@@
The embeddings for these anchors are computed from the text embeddings of the sensor readings using a pre-trained CLIP text embedding model. Subsequently, these anchor embeddings are blended to generate embeddings for new sensor readings. As illustrated in Figure 6(a), for a single sensor reading, such as temperature, the embedding can be derived through linear interpolation between the corresponding anchors. For scenarios involving more than two anchors (e.g., with multiple representative points), the embedding is calculated using linear interpolation between the neighboring anchors (Figure 6(b)). Key challenges arise from the increasing number of sensors and anchors and the complexity of handling their combinations:
@@
We present a showcase example with humidity sensors to demonstrate the effectiveness of our method in visual space. We generate images using Stable Diffusion with embeddings computed by CLIP, based on the prompt template: ”Urban skyline with buildings under $x$ AQI, where $x$ can substitute with different values. We compare the results of our embedding interpolation method (referred to as the ‘interpolated’ set) with a baseline approach (referred to as the ‘direct’ set), where embeddings were directly fed into the diffusion model. The comparative results are shown in Figure 7. For interpolation, we use embeddings generated at the minimum and maximum values of $x$.
@@
As Air Quality Index (AQI) values increase, we expect to see a continuous progression of visual characteristics. These include a gradual transition in atmospheric coloration from clear blue to yellow-gray and finally to brown-red hues, as well as an increasing prominence of light source diffraction phenomena, resulting in the formation of halos or luminous effects through particulate matter. While this progression is evident in the interpolated set, we can clearly see this inconsistence in the direct set. Specifically, the image corresponding to AQI of 90 (the second image from the left) in the direct set fails to align with the expected progression.
To quantify these differences, we computed the cosine similarity of each embedding relative to the embeddings of the two extreme conditions (AQI of 44 and 314). This analysis yielded two sequences from the direct set, shown as blue lines in the top two plots in Figure 7. The corresponding sequences from the interpolated set are shown as orange lines in the same plots. To further evaluate the disruption in monotonic progression, we employed Kendall’s Tau coefficient, which measures the degree of monotonicity, with a value approaching 1 indicates a perfect monotonic increase, while a value approaching -1 indicates a perfect monotonic decrease. It is defined as $\tau=\frac{2(P-Q)}{n(n-1)}$, where $P$ is the number of concordant pairs, $Q$ is the number of discordant pairs, and $n$ is the total number of pairs.
The results indicate that the embedding interpolation method achieves perfect monotonicity, as reflected by Kendall’s Tau values and the smooth progression of the orange curves. This contrasts with the inconsistencies observed in the direct set, as shown in the valley of the blue curve.
This example demonstrates that the embedding interpolation approach provides a more consistent and accurate representation of sensor data in the sensor visual space. By interpolation embeddings from anchor points, we achieve smooth transitions in the generated images corresponding to incremental changes in sensor readings. Next, we extend this design to cases with more than one sensor reading.